OpenStudy (anonymous):

Write the first 5 terms of the following sequences a =5n - 3 ^n

OpenStudy (anonymous):

Just plug in n from 1 to 5, so ex. 8th term would be 5(8)-3 = 40-3=37, except plug in 1,2,3,4 and 5